| ID | 
Date | 
Author | 
Category | 
Subject | 
Text | 
  | 
| 
  50  
 | 
Tue Apr  1 20:37:04 2025 | 
Jeroen | DAQ | Change of settings for shorter acq. window | Acquisition stopped at 20:36. 
 
 
First file from this setting: RSA01-2025.04.01.03.52.59.247.tiq
 | 
   | 
| 
  49  
 | 
Tue Apr  1 20:19:23 2025 | 
Jeroen, Zac, Esther | Accelerator | Beam fell out | > > No beam from 19:57:41.
 
> 
 
> Beam back from 20:02 | 
    | 
| 
  48  
 | 
Tue Apr  1 20:02:19 2025 | 
Jeroen, Zac, Esther | Accelerator | Beam fell out | > No beam from 19:57:41.
 
 
Beam back from 20:02 | 
   | 
| 
  47  
 | 
Tue Apr  1 20:00:14 2025 | 
Jeroen, Zac, Esther | Accelerator | Beam fell out | No beam from 19:57:41. | 
   | 
| 
  46  
 | 
Tue Apr  1 17:08:13 2025 | 
JG, ID, MC, CG | Detectors | pleiades in position | The pleiades detectors are now moved into 
a position which does not disturb the stored 
beam. 
 | 
   | 
| 
  45  
 | 
Tue Apr  1 16:47:32 2025 | 
Ruijiu, Jan, Shahab | Analysis | The frequency of 194Pb82+ is changed to 408.600 MHz |   | 
     | 
| 
  44  
 | 
Tue Apr  1 16:44:32 2025 | 
Jeroen, Esther | General | No SIS beam since 14:35 | Beam resumed at 16:40.  | 
     | 
| 
  43  
 | 
Tue Apr  1 15:38:50 2025 | 
ruijiu chen | Analysis | frequency drift |   | 
     | 
| 
  42  
 | 
Tue Apr  1 15:32:59 2025 | 
Zac | DAQ | NTCAP Restarted | NTCAP Has been restarted (we restart every 
day around 15:00 so data goes in a new folder) | 
   | 
| 
  41  
 | 
Tue Apr  1 14:54:24 2025 | 
all | General | No SIS beam since 14:35 |   | 
   | 
| 
  40  
 | 
Tue Apr  1 14:25:21 2025 | 
ruijiu  | Analysis | The frequency of 194Pb82+ has been changed from 408.606MHz to 408.602MHy during last 12 hours | The following images display the frequency 
spectrum at the start of every hour from 
3:00 to 14:00. | 
 12x   | 
| 
  39  
 | 
Tue Apr  1 14:23:01 2025 | 
JG | General | Pleiades pocket position tests | We started playing with the detector pockets 
of the pleiades arrays at 14:23 01.04.2025.
 
 | 
   | 
| 
  38  
 | 
Tue Apr  1 14:17:07 2025 | 
Ruijiu | Analysis | PID of 194Pb82+ 2025.04.01.14.13.07 | datafile = "/data.local/G22-00018_00203/results/analyzers/RSA01/RSA01-2025.04.01.14.13.07.068.tiq_spectrum.npz"
 
filep = "/data.local/G22-00018_00203/scripts/OnlineDataAnalysisSystem/rionid_exercises/exercise_194Pb_Exp_2025/New_TEline2022-ESR-194Pb-208Pb-nostripper_524MeV.lpp"
 
alphap = "1.399"
 | 
    | 
| 
  37  
 | 
Tue Apr  1 14:02:23 2025 | 
Jelena B, Siddharth | General | Lassie |   | 
    | 
| 
  36  
 | 
Tue Apr  1 12:58:27 2025 | 
Jelena B, Siddharth | General | Lassie |   | 
    | 
| 
  35  
 | 
Tue Apr  1 11:47:14 2025 | 
Jelena B, Siddharth | General | Lassie 2025-04-01_11-46-35-517_tcl1064.png  |   | 
    | 
| 
  34  
 | 
Tue Apr  1 10:34:20 2025 | 
Jelena B, Siddharth | General | Resolved issue with SIS18 | Beam is back with some fluctuations.
 
We did not have beam from 9:34 until 10:08, 
around 35 minutes.  | 
    | 
| 
  33  
 | 
Tue Apr  1 10:03:56 2025 | 
Jelena B, Siddharth | General | Technical issue with SIS18  | From HKR info: there is a technical issue 
with SIS18 and we will have no beam for next 
25 minutes.  | 
    | 
| 
  32  
 | 
Tue Apr  1 08:36:25 2025 | 
Jelena B, Siddharth | General | Lassie 2025-04-01_08-37-32-155_tcl1064.png |   | 
    | 
| 
  31  
 | 
Tue Apr  1 06:53:51 2025 | 
Jelena B,Siddharth | General | Lassie Monitor |   | 
    |